# Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback

### v1

@@ 3 The Proposed Algorithm: RL-LOW
In this section, we describe our computationally and statistically efficient algorithm for offline RL with preference feedback based on the novel idea of locally optimal weights for estimating the relative reward of each pair of actions. This algorithm, called RL-LOW, is simple and is presented formally in Algorithm 1. Before we describe its components, we introduce some notations.
First, we denote $B_{k,i,j}$ as the empirical success rate with the comparison of actions $i$ and $j$, i.e., for $k\in\mathcal{S}$ and $i,j\in\mathcal{A}$ with $N_{k,i,j}>0$
$B_{k,i,j}\coloneqq\frac{1}{nN_{k,i,j}}\sum_{\iota=1}^{n}\sigma_{\iota}\mathbbm{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\},$ (5)
and $B_{k,j,i}\coloneqq 1-B_{k,i,j}$. If $N_{k,i,j}=N_{k,j,i}=0$, we define $B_{k,i,j}=B_{k,j,i}=0$. Subsequently, certain empirical success rates may exhibit magnitudes that are either excessively large or small. We clip them by means of the following operation:
${B}_{k,i,j}^{\rm CLP}=\mathrm{CLIP}_{L}({B}_{k,i,j})\quad\mbox{where}\quad\mathrm{CLIP}_{L}(a)=\left\{\begin{array}[]{cc}\frac{\exp(2L)}{1+\exp(2L)}&a>\frac{\exp(2L)}{1+\exp(2L)}\\ \frac{1}{1+\exp(2L)}&a<\frac{1}{1+\exp(2L)}\\ a&\text{otherwise}\end{array}\right..$ (6)
In accordance with Assumption 2.1, the implicit rewards are bounded by $L$. Consequently, within our BTL model framework, the success rate of each comparison necessarily falls within the interval $\big{[}\frac{1}{1+\exp(2L)},\frac{\exp(2L)}{1+\exp(2L)}\big{]}$. We exploit this in (6) to ensure that the implementation of our clip operation is consistent with the model’s dynamics. We are now ready to introduce the notion of locally optimal weights, which plays a central role in the estimation of the rewards.
@@
0:  Dataset $\mathcal{D}=\{(s_{i},a^{(1)}_{i},a^{(2)}_{i},\sigma_{i})\}_{i=1}^{n}$ and feature maps $\phi=\{\phi(k,i)\}_{k\in\mathcal{S},i\in\mathcal{A}}$
0:  The estimated best action $\hat{i}_{k}\in\mathcal{A}$ of each state $k\in\mathcal{S}$.
1:  Compute the sample proportions $N_{k,i,j}\leftarrow\frac{1}{n}\sum_{\iota=1}^{n}\mathbbm{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}$.
2:  For $k\in\mathcal{S}$ and $i,j\in\mathcal{A}$ such that $N_{k,i,j}>0$, compute the success rate of each pair of comparisons using Eqn. (5).
3:  Clip the success rate using the knowledge of $L$ and using Eqn. (6).
@@
(Locally Optimal Weight) For an consistent instance $v$, let $\mathcal{U}_{k,i,j}=\{u\in\mathbb{R}^{S\times A\times A}:\phi(k,i)-\phi(k,j)=\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}}u_{k^{\prime},i^{\prime},j^{\prime}}(\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}))\;\mbox{and}\;u_{k^{\prime},i^{\prime},j^{\prime}}=0\;\mbox{if}\;N_{k^{\prime},i^{\prime},j^{\prime}}=0\}$. We say that $w^{(k,i,j)}=(w_{k^{\prime},i^{\prime},j^{\prime}}^{(k,i,j)})_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}}$ is a set of locally optimal weights for $(k,i,j)\in\mathcal{S}\times\mathcal{A}^{2}$ with $i\neq j$ if
@@
The weights in Eqn. (7) are described as “locally optimal” because they are customized to each $(k,i,j)$ tuple. Hence, $w^{(k,i,j)}$ is local to $(k,i,j)$. This is a novelty in the design of our algorithm.
By the definition of the consistency of an instance (cf. Definition 2.2), there exists a subset $\beta\subset\mathcal{S}\times\mathcal{A}^{2}$ such that $\phi(k,i)-\phi(k,j)\in{\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\beta\}.$ Hence, there exists a locally optimal weight for every pair of actions given a consistent instance. In addition, $w^{(k,i,j)}$ can be calculated efficiently by its analytic form (see details in Appendix C).
Equipped with the definition of locally-optimal weights, we now provide an estimate of the relative reward for state $k\in\mathcal{S}$ and pair of action $(i,j)\in\mathcal{A}^{2}$ with $i\neq j$ as follows:
$\hat{r}_{k,i,j}=\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}}w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\,\log\Big{(}\frac{B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}{1-B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}\Big{)}.$ (8)
The term $\frac{(u_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}$ in Eqn. (7) is a proxy for the variance introduced by the pair of actions $(i^{\prime},j^{\prime})$ in state $k^{\prime}$ when associated with the coefficient $u_{k^{\prime},i^{\prime},j^{\prime}}$ in the linear combination of the definition of $\mathcal{U}_{k,i,j}$. Our objective is to minimize the cumulative variance proxy for $(k,i,j)$, thus enhancing the precision of the estimate of the relative reward for $(k,i,j)$ for the purposes of establishing the tightest possible concentration result for subGaussian random variables (see Appendix E).
Finally, for any $k\in\mathcal{S}$, let $\hat{i}_{k}\in\hat{\mathcal{I}}_{k}:=\{i\in\mathcal{A}:\hat{r}_{k,i,j}\geq 0,\forall\,j\neq i\}$ be any estimate of the best action under state $k$. It is natural to wonder whether $\hat{i}_{k}$ exists, i.e., whether the set ^$\hat{\mathcal{I}}_{k}$ is empty. The following proposition answers this in the affirmative.
@@
For any consistent instance $v$ and using estimate of the best action under each state $k$ as prescribed by RL-LOW, we have |^$\lvert\hat{\mathcal{I}}_{k}\rvert\geq 1$, i.e., $\hat{i}_{k}$ exists.
@@ 3.1 Upper Bound of RL-LOW
In this section, we provide an instance-dependent upper bound of the simple regret for Rl-LOW. In addition, we also provide a worst-case upper bound as a by-product. First, we define an instance-dependent hardness parameter $H(v)$. Let
$H(v)\coloneqq\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}}\ \frac{\gamma_{k,i}}{\Delta_{k,i}^{2}}\quad\mbox{where}\quad\gamma_{k,i}\coloneqq\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{(w^{(k,i,i_{k}^{*})}_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}.$ (9)
The parameter $\gamma_{k,i}$ exhibits a positive correlation with the variance proxy of the relative empirical reward $\hat{r}_{k,i,i_{k}^{*}}$. Consequently, the ratio $\frac{\gamma_{k,i}}{\Delta_{k,i}^{2}}$ in the definition of $H(v)$ serves as a quantitative measure of the difficulty that the empirical reward of a suboptimal action $i$ surpasses that of $i^{*}_{k}$ in state $k$.
@@
From Theorem 3.3, it is evident that the upper bound decays exponentially fast and the exponent is a function of an instance-dependent hardness term $H(v)$. This is the first instance-dependent analysis in offline reinforcement learning with preference feedback.
It is natural to wonder why we do not devise an instance-dependent analysis of or modification to Pessimistic MLE which was developed by Zhu et al., (2023). Note that Pessimistic MLE is designed to perform well with high probability and not necessarily in expectation. In particular, the regret bound of Pessimistic MLE holds with probability at least $1-\delta$. Hence, to ensure the regret is less than $\exp(-\Omega(n/H(v)))$, one should set the failure probability $\delta$ to be $\exp(-\Theta(n/H(v)))$, which is not possible as $H(v)$ is unknown to the algorithm (since $\theta$ is also unknown).
We further provide a worst-case upper bound for RL-LOW as follows.
@@
$\!\mathbb{E}^{\rm RL\mbox{-}LOW}_{v}\left[R_{n}\right]\!\leq\!\frac{\sum\limits_{\begin{subarray}{c}k,i:i\neq i^{*}_{k}\end{subarray}}\rho_{k}(\sqrt{\gamma_{k,i}}\!+\!\tilde{\gamma}_{k,i})}{C_{\rm wup}\sqrt{n}}\;\,\mbox{where}\;\,\tilde{\gamma}_{k,i}\!=\!\sum_{\begin{subarray}{c}k^{\prime},i^{\prime},j^{\prime}:N_{k^{\prime},i^{\prime},j^{\prime}}>0\end{subarray}}\frac{\lvert w^{(k,i,i_{k}^{*})}_{k^{\prime},i^{\prime},j^{\prime}}\rvert}{\sqrt{N_{k^{\prime},i^{\prime},j^{\prime}}}}$ (11)
@@
We note that in Zhu et al., (2023), the high probability upper bound is of the form $O\big{(}\sqrt{n^{-1}\log(1/\delta)}\big{)}$. Hence, if we desire a bound in expectation, we obtain, through the law of total probability, a bound of the form $\mathbb{E}[R_{n}]=O\big{(}\sqrt{n^{-1}\log(1/\delta)}+\delta\big{)}$. Minimizing this bound over $\delta$ yields $\mathbb{E}[R_{n}]=O\big{(}\sqrt{n^{-1}\log n}\big{)}$, which is slightly inferior to our upper bound $O(n^{-1/2})$.
@@ 4 Instance-Dependent Lower Bound
In this section, we derive the first-of-its-kind instance-dependent lower bound on offline RL with preference feedback. Before we present our bound, we present some auxiliary lemmas that are potentially instrumental in deriving lower bounds on other preference-based RL problems.
Given any instance $v$, we let $P_{v}^{(n)}$ denote the joint distribution of the associated labels $\{\sigma_{i}\}_{i=1}^{n}$. The following lemma provides an estimate of the Kullback–Leibler (KL) divergence between instances $v$ and $v^{\prime}$ that share the same parameters except for the latent vector $\theta$ that defines the reward in (1).
@@
where
This lemma demonstrates that when the rewards are bounded, the weighted sum of squared differences of the relative rewards can be used to approximate the KL divergence between the distributions of two instances. The approximation is precisely $\tilde{D}(v,v^{\prime})$ defined in (13). Furthermore, for any $\mathbb{z}\in\mathbb{R}^{d}$, $\eta\in\mathbb{R}$ and consistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$, we let ${\rm Alt}(v,\mathbb{z},\eta)$ be the set of instances that share the same instance parameters except for $\theta$ and satisfies $\langle\mathbb{z},\theta^{\prime}-\theta\rangle=\eta$ for all $v^{\prime}=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta^{\prime})\in{\rm Alt}(v,\mathbb{z},\eta)$. The following lemma states a useful property that relates the ${\rm Alt}$ set to the “approximate KL divergence” $\tilde{D}$.
@@
Let $\mathcal{G}$ be an arbitrary orthonormal basis of ${\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}\text{ and }N_{k^{\prime},i^{\prime},j^{\prime}}>0\}$. Also let $[\mathbf{w}]_{\mathcal{G}}$ denote the column vector that represents $\mathbf{w}$ under basis $\mathcal{G}$ (Meyer,, 2000, Chapter 4). Define the matrix
@@
In addition, let $(\bar{i}$, $\bar{k})$ be the state-action pair that attains maximum in the definition of hardness in Eqn. (9). Define the subset of instances
$\mathcal{Q}=\bigg{\{}v\mbox{ consistent}:\frac{\gamma_{\bar{k},\bar{i}}}{\Delta_{\bar{k},\bar{i}}^{2}}\geq 4\frac{\gamma_{k,i}}{\Delta_{k,i}^{2}}\;\;\forall\,(k,i)\neq(\bar{k},\bar{i})\;\mbox{and}\;i\neq i^{*}_{k}\bigg{\}}.$ (16)
@@
Theorem 4.3 is an instance-dependent lower bound for all instances in the set $\mathcal{Q}$. The condition that defines $\mathcal{Q}$ in Eqn. (16) ensures that the hardness quantities $H(v)$ and $H(v^{\prime})$ have the same order. Since instances in $\mathcal{Q}$ cover all possible hardness values $H(v)$ (i.e., for every hardness values $h>0$, there exists an instance in $\mathcal{Q}$ of hardness $h$), we conclude that for any consistent instance $v$ and (small) $\epsilon\in(0,1)$, there does not exist any policy $\pi$ that achieves

### v2

@@ 3 The Proposed Algorithm: RL-LOW
In this section, we describe our computationally and statistically efficient algorithm for offline RLHF with pairwise comparisons based on the novel idea of locally optimal weights for estimating the relative reward of each pair of actions. For clarity in exposition, this section is devoted to the setting of a static state distribution $\rho$; this is done as a foundational step before we extend the ideas to the MDP setting of RL in Section 5.
Our proposed algorithm, called RL-LOW, is simple and is presented formally in Algorithm 1. Before we describe its components, we introduce some notations.
Let $B_{k,i,j}$ be the empirical success rate with the comparison of $i$ and $j$, i.e., for $k\in\mathcal{S}$ and $i,j\in\mathcal{A}$ with $N_{k,i,j}>0,$
$B_{k,i,j}\coloneqq\frac{1}{nN_{k,i,j}}\sum_{\iota=1}^{n}\sigma_{\iota}\mathds{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\},$ (5)
and $B_{k,j,i}\coloneqq 1-B_{k,i,j}$. If $N_{k,i,j}=N_{k,j,i}=0$, we define $B_{k,i,j}=B_{k,j,i}=0$. Subsequently, certain empirical success rates may exhibit magnitudes that are either excessively large or small. We clip them by means of the following operation: ${B}_{k,i,j}^{\rm CLP}=\mathrm{CLIP}_{L}({B}_{k,i,j})$, where
[…]
Per Assumption 2.1, the implicit rewards are bounded by $L$. Consequently, within our BTL model framework, the success rate of each comparison necessarily falls within the interval $\big[\frac{1}{1+\exp(2L)},\frac{\exp(2L)}{1+\exp(2L)}\big]$. We exploit this in Eqn. (6) to ensure that the implementation of our clip operation is consistent with the model’s dynamics. We are now ready to introduce the notion of locally optimal weights, which plays a central role in the estimation of the rewards.
@@
0: Dataset $\mathcal{D}=\{(s_{i},a^{(1)}_{i},a^{(2)}_{i},\sigma_{i})\}_{i=1}^{n}$ and feature maps $\phi=\{\phi(k,i)\}_{k\in\mathcal{S},i\in\mathcal{A}}$ .
0: The estimated best action $\hat{i}_{k}\in\mathcal{A}$ for each $k\in\mathcal{S}$.
1: Compute the sample proportions $N_{k,i,j}\leftarrow\frac{1}{n}\sum_{\iota=1}^{n}\mathds{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}$.
2: For $k\in\mathcal{S}$ and $i,j\in\mathcal{A}$ such that $N_{k,i,j}>0$, compute the success rate $B_{k,i,j}$ using Eqn. (5).
3: Compute ${B}_{k,i,j}^{\rm CLP}$ by clipping $B_{k,i,j}$ through Eqn. (6).
@@
(Locally Optimal Weights) For an consistent instance $v$, let $\mathcal{U}_{k,i,j}=\{u\in\mathbb{R}^{S\times A\times A}:\phi(k,i)-\phi(k,j)=\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}}u_{k^{\prime},i^{\prime},j^{\prime}}(\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}))\;\mbox{and}\;u_{k^{\prime},i^{\prime},j^{\prime}}=0\;\mbox{if}\;N_{k^{\prime},i^{\prime},j^{\prime}}=0\}$. We say that $w^{(k,i,j)}=(w_{k^{\prime},i^{\prime},j^{\prime}}^{(k,i,j)})_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}}$ is a set of locally optimal weights for $(k,i,j)\in\mathcal{S}\times\mathcal{A}^{2}$ with $i\neq j$ if
@@
The weights in Eqn. (7) are described as “locally optimal” because they are specifically customized to each $(k,i,j)$ tuple. Hence, $w^{(k,i,j)}$ is local to $(k,i,j)$. This is a novelty in the design of our algorithm.
By the definition of the consistency of an instance (cf. Definition 2.2), there exists a subset $\beta\subset\mathcal{S}\times\mathcal{A}^{2}$ such that $\phi(k,i)-\phi(k,j)\in{\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\beta\}.$ Hence, there exists a locally optimal weight for every pair of actions given a consistent instance. In addition, $w^{(k,i,j)}$ can be calculated efficiently by its analytic form (see details in Appendix F [^3] Following AAAI policy, the appendices are provided on the arxiv (Chen and Tan 2025).).
Equipped with the definition of locally optimal weights, we now provide an estimate of the relative reward for state $k\in\mathcal{S}$ and pair of action $(i,j)\in\mathcal{A}^{2}$ with $i\neq j$ as follows:
$\hat{r}_{k,i,j}=\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}}w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\,\log\Big(\frac{B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}{1-B^{\rm CLP}_{k^{\prime},i^{\prime},j^{\prime}}}\Big),$ (8)
and we define $\hat{r}_{k,i,i}=0$ for all $k\in\mathcal{S}$ and $i\in\mathcal{A}$. The term $\frac{(u_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}$ in Eqn. (7) is a proxy for the variance introduced by the pair of actions $(i^{\prime},j^{\prime})$ in state $k^{\prime}$ when associated with the coefficient $u_{k^{\prime},i^{\prime},j^{\prime}}$ in the linear combination of the definition of $\mathcal{U}_{k,i,j}$. Our objective is to minimize the cumulative variance proxy for $(k,i,j)$, thus enhancing the precision of the estimate of the relative reward for $(k,i,j)$ for the purposes of establishing the tightest possible concentration result for subGaussian random variables (see Appendix H).
Finally, for any $k\in\mathcal{S}$, let $\hat{i}_{k}\in\hat{\mathcal{I}}_{k}:=\{i\in\mathcal{A}:\hat{r}_{k,i,j}\geq 0,\forall\,j\neq i\}$ be any estimate of the best action under state $k$. It is natural to wonder whether $\hat{i}_{k}$ exists, i.e., whether the set $\hat{\mathcal{I}}_{k}$ is empty. The following proposition answers this in the affirmative.
@@
For any consistent instance $v$ and using estimate of the best action $\hat{i}_{k}$ under each state $k$ as prescribed by RL-LOW, we have $\lvert\hat{\mathcal{I}}_{k}\rvert\geq 1$ and
$\operatorname*{argmax}_{i\in\mathcal{A}}\hat{r}_{k,i,j_{1}}=\operatorname*{argmax}_{i\in\mathcal{A}}\hat{r}_{k,i,j_{2}}=\hat{\mathcal{I}}_{k}\quad\mbox{for any}\ j_{1},j_{2}\in\mathcal{A}.$
Computational Complexity:
Proposition 3.2 obviates the need to compute all values of $\hat{r}_{k,i,j}$ for each $(k,i,j)\in\mathcal{S}\times\mathcal{A}^{2}$. We demonstrate that the RL-LOW algorithm can be efficiently implemented with a computational complexity of $\mathcal{O}(SAd+nd^{2}+d^{3})$, as the term $SAd$ corresponds to the natural process of scanning the feature vectors for all state-action pairs. The terms $nd^{2}+d^{3}$ are typical in scenarios involving a linear reward structure, such as in linear regression. It is worth noting that the term $SAd$ can be removed if we do not need to output $\hat{i}_{k}$ for each $k\in\mathcal{S}$, but rather a parametric function $\hat{i}(k;\vartheta)$ is to be learned, and the overall computational complexity becomes $\mathcal{O}(nd^{2}+d^{3})$; see details in Appendix F.
@@ 3.1 Upper Bound of RL-LOW
In this section, we provide an instance-dependent upper bound of the simple regret for Rl-LOW, and we also provide a worst-case upper bound as a by-product. First, we define an instance-dependent hardness parameter $H(v)$. Let
$H(v)\coloneqq\max_{k\in\mathcal{S},i\in\mathcal{A}:i\neq i^{*}_{k}}\ \frac{\gamma_{k,i}}{\Delta_{k,i}^{2}},$ (9)
$\mbox{where }\gamma_{k,i}\coloneqq\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{(w^{(k,i,i_{k}^{*})}_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}.$
The parameter $\gamma_{k,i}$ exhibits a positive correlation with the variance proxy of the relative empirical reward $\hat{r}_{k,i,i_{k}^{*}}$. Consequently, the ratio $\frac{\gamma_{k,i}}{\Delta_{k,i}^{2}}$ in the definition of $H(v)$ serves as a quantitative measure of the difficulty that the empirical reward of a suboptimal action $i$ surpasses that of $i^{*}_{k}$ in state $k$; see more intuitive explanations in Appendix C.
@@
From Theorem 3.3, it is evident that the upper bound decays exponentially fast and the exponent is a function of an instance-dependent hardness term $H(v)$. This is the first instance-dependent analysis in offline reinforcement learning with pairwise comparisons.
It is natural to wonder why we do not devise an instance-dependent analysis of or modification to the existing Pessimistic MLE in Zhu et al. (2023). Note that Pessimistic MLE is designed to perform well with high probability and not necessarily in expectation. In particular, the regret bound of Pessimistic MLE holds with probability at least $1-\delta$. Hence, to ensure the regret is less than $\exp(-\Omega(n/H(v)))$, one should set the failure probability $\delta$ to be $\exp(-\Theta(n/H(v)))$, which is not possible as $H(v)$ is unknown to the algorithm (since $\theta$ is also unknown). We further provide a worst-case upper bound for RL-LOW as follows.
@@
$\mathbb{E}^{\rm RL\mbox{-}LOW}_{v}\left[R_{n}\right]\!\leq\!\frac{\sum\limits_{\begin{subarray}{c}k,i:i\neq i^{*}_{k}\end{subarray}}\rho_{k}(\sqrt{\gamma_{k,i}}\!+\!\tilde{\gamma}_{k,i})}{C_{\rm wup}\sqrt{n}}\;\,$ (11)
$\mbox{where}\;\,\tilde{\gamma}_{k,i}\!=\!\sum_{\begin{subarray}{c}k^{\prime},i^{\prime},j^{\prime}:N_{k^{\prime},i^{\prime},j^{\prime}}>0\end{subarray}}\frac{\lvert w^{(k,i,i_{k}^{*})}_{k^{\prime},i^{\prime},j^{\prime}}\rvert}{\sqrt{N_{k^{\prime},i^{\prime},j^{\prime}}}}$
@@
We note that in Zhu et al. (2023), the high probability upper bound is of the form $O\big(\sqrt{n^{-1}\log(1/\delta)}\big)$ for the dependency of $n$ and $\delta$. Hence, if we desire a bound in expectation, we obtain, through the law of total probability, a bound of the form $\mathbb{E}[R_{n}]=O\big(\sqrt{n^{-1}\log(1/\delta)}+\delta\big)$. Minimizing this bound over $\delta$ yields $\mathbb{E}[R_{n}]=O\big(\sqrt{n^{-1}\log n}\big)$. In terms of the dependence on $n$, it exhibits a performance that is slightly inferior to our established upper bound $O(n^{-1/2})$ of RL-LOW.
@@ 4 Instance-Dependent Lower Bound
In this section, we derive the first-of-its-kind instance-dependent lower bound on offline RLHF with pairwise comparisons. Before we present our bound, we present some auxiliary lemmas that are potentially instrumental in deriving lower bounds on other preference-based RL problems.
Given any instance $v$, we let $P_{v}^{(n)}$ denote the joint distribution of the associated labels $\{\sigma_{i}\}_{i=1}^{n}$. The following lemma provides an estimate of the Kullback–Leibler (KL) divergence between instances $v$ and $v^{\prime}$ that share the same parameters except for the latent vector $\theta$ as in Eqn. (1).
@@
$2n\exp(-4R_{\max})\leq\frac{D_{\rm KL}(P_{v}^{(n)}\|P_{v^{\prime}}^{(n)})}{\tilde{D}(v,v^{\prime})}\leq 2n\exp(2R_{\max})$
$\text{where }\tilde{D}(v,v^{\prime})=\sum_{k\in\mathcal{S},i,j\in\mathcal{A}}N_{k,i,j}(\langle\phi(k,i)-\phi(k,j),\theta-\theta^{\prime}\rangle)^{2},$ (12)
and $R_{\max}=\max_{k\in\mathcal{S},i\in\mathcal{A}}\max\{\lvert\langle\phi(k,i),\theta\rangle\rvert,\lvert\langle\phi(k,i),\theta^{\prime}\rangle\rvert\}$ is the maximum absolute reward in these two instances.
This lemma demonstrates that when the rewards are bounded, the weighted sum of squared differences of the relative rewards can be used to approximate the KL divergence between the distributions of two instances. The approximation is precisely $\tilde{D}(v,v^{\prime})$ defined in Eqn. (12). Furthermore, for any $\mathbb{z}\in\mathbb{R}^{d}$, $\eta\in\mathbb{R}$ and consistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$, we let ${\rm Alt}(v,\mathbb{z},\eta)$ be the set of instances that share the same instance parameters except for $\theta$ and satisfies $\langle\mathbb{z},\theta^{\prime}-\theta\rangle=\eta$ for all $v^{\prime}=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta^{\prime})\in{\rm Alt}(v,\mathbb{z},\eta)$. The following lemma states a useful property that relates the $\rm Alt$} set to the “approximate KL divergence” $\tilde{D}$.
@@
Let $\mathcal{G}$ be an arbitrary orthonormal basis of ${\rm Span}\{\phi(k^{\prime},i^{\prime})-\phi(k^{\prime},j^{\prime}):(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}\text{ and }N_{k^{\prime},i^{\prime},j^{\prime}}>0\}$. Also let $\mathbf{w}]_{\mathcal{G}}$ denote the column vector that represents $\mathbf{w}$ under basis $\mathcal{G}$ (Meyer 2000, Chapter 4). Define the matrix
@@
In addition, let $(\bar{i}$, $\bar{k})$ be the state-action pair that attains maximum in the definition of hardness in Eqn. (3.1). Define the subset of instances
$\mathcal{Q}=\bigg\{v\mbox{ consistent}:\frac{\gamma_{\bar{k},\bar{i}}}{\Delta_{\bar{k},\bar{i}}^{2}}\geq\frac{4\gamma_{k,i}}{\Delta_{k,i}^{2}}\;\;\forall\,(k,i)\neq(\bar{k},\bar{i})\;\mbox{,}\;i\neq i^{*}_{k}\bigg\}.$ (15)
@@
Theorem 4.3 is an instance-dependent lower bound for all instances in the set $\mathcal{Q}$. The condition that defines $\mathcal{Q}$ in Eqn. (15) ensures that the hardness quantities $H(v)$ and $H(v^{\prime})$ have the same order. Since instances in $\mathcal{Q}$ cover all possible hardness values $H(v)$ (i.e., for every hardness values $h>0$, there exists an instance in $\mathcal{Q}$ of hardness $h$), we conclude that for any (small) $\epsilon\in(0,1)$, there does not exist any algorithm $\Pi$ that achieves for all consistent instance $v$,
@@
5 Extension to the MDP Setting of RL
Similar to Zhu et al. (2023, Section 1), our definition of simple regret is based on the static state distribution $\rho$ in the previous sections. In this section, we extend our results to the MDP setting of RL when the transition probabilities $P(k^{\prime}|k,i)$ for $(k^{\prime},k,i)\in\mathcal{S}^{2}\times\mathcal{A}$ are assumed to be known. This assumption is consistent with  Zhu et al. (2023, Section 5), and we also provide a motivational example for this assumption in Appendix A. Given the transition probabilities $P(k^{\prime}|k,i)$ and an MDP policy $\pi$, we let $d^{\pi}$ denote the state distribution (Sutton and Barto 2018, Section 9.2) under $\pi$. Without loss of generality, we assume the MDP policies are deterministic, and we denote $\pi(k)\in\mathcal{A}$ to be the output action of $\pi$ under state $k$. Let $\pi^{*}$ denote the optimal MDP policy that is assumed to be unique, i.e.,
$\pi^{*}=\operatorname*{argmax}_{\pi}\mathbb{E}_{k\sim d^{\pi}}\big[r_{k,\pi(k)}\big].$
Then, we define the simple regret of any MDP policy $\pi$ as
$R^{\rm MDP}(\pi)=\mathbb{E}_{k\sim d^{\pi^{*}}}\big[r_{k,\pi^{*}(k)}]-\mathbb{E}_{k\sim d^{\pi}}\big[r_{k,\pi(k)}\big].$ (17)
We now adapt our RL-LOW to the MDP setting by redefining the output as an MDP policy:
$\hat{\pi}_{\rm out}\in\operatorname*{argmax}_{\pi}\mathbb{E}_{k\sim d^{\pi}}[\hat{r}_{k,\pi(k),j^{\dagger}}],$
where $j^{\dagger}\in\mathcal{A}$ is arbitrarily fixed (e.g., $j^{\dagger}=1$). We simply call this adaptation RL-LOW-MDP. The upper bound on its simple regret is stated as follows.
Theorem 5.1.
(Instance-Dependent Upper Bound for RL-LOW-MDP) Given any consistent instance $v$, for all sufficiently large $n$,
$\mathbb{E}^{\rm RL\mbox{-}LOW\mbox{-}MDP}_{v}\left[R^{\rm MDP}(\hat{\pi}_{\rm out})\right]\leq\exp\bigg(-\frac{n}{C_{\rm MDP}\cdot H_{\rm MDP}(v)}\bigg)$
where $C_{\rm MDP}>0$ is a universal constant,
$H_{\rm MDP}(v)\coloneqq\max_{\pi\neq\pi^{*}}\frac{\gamma^{\rm MDP}(\pi)}{(\mathbb{E}_{k\sim d^{\pi^{*}}}\big[r_{k,\pi^{*}(k)}]-\mathbb{E}_{k\sim d^{\pi}}\big[r_{k,\pi(k)}\big])^{2}},$
$\gamma^{\rm MDP}(\pi)\coloneqq\max_{k:\pi(k)\neq\pi^{*}(k)}\ \sum_{k^{\prime},i^{\prime},j^{\prime}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{(w^{(k,\pi(k),\pi^{*}(k)}_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}.$
In the presence of the MDP, $H_{\rm MDP}(v)$, which is a generalization of $H(v)$ in Eqn. (3.1), turns out to be the instance-dependence hardness parameter of the problem. The proof of Theorem 5.1 is provided in Appendix H. It is important to observe that there exist MDPs (e.g., $P(k|k,i)=1$ or $S=1$) such that Theorem 5.1 particularizes to Theorem 3.3. Moreover, the lower bound in Theorem 4.3 is also applicable to the present more general MDP setting when the transition probability kernel $P(k^{\prime}|k,i)$ is independent of $(k,i)$ and $k^{\prime}$ follows the distribution $\rho$. Admittedly, the complexity of the problem increases substantially when the transition probabilities are unknown; this aspect warrants further investigation in future studies. Our findings serve as an initial step in exploring instance-dependent bounds in offline RLHF.